OpenStudy (anonymous):

find the sum of 3/7 + 1/2

OpenStudy (anonymous):

13/14

OpenStudy (anonymous):

do you need me to explain how to add them?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Find a common denominator, and then combine the top of the fractions with addition. In this case, 14 is the lowest common denominator between 7 and 2, so -- 1/2=7/14 & 3/7 = 6/14 7/14 + 6/14 = 13/14

OpenStudy (anonymous):

you need to find the LCD (least common denominator) you do that by seeing what the lowest number but the denominators(bottom number of the fraction) can go into. 7 and 2 are the denominators. 2 cannot go evenly into 7. 7+7=14. 7 can go evenly into 14 as well as 2. now you have to switch both denominators to 14. the numerators change too. 7 goes into 14, 2 times. 2 times 3 = 6. new fraction is 6/14. next, 2 goes into 14, 7 times. 7 times 1 = 7. new fraction is 7/14. add the numerators. 6+7=13. do not add the denominators. they stay the same. therefor, 13/14 is the sum. :) hope this helps